OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (-6, 3). Answer y = 3/2x - 1 y = 3/2x + 3 y = - 2/3x - 1 y = - 2/3x + 3

OpenStudy (angela210793):

perpendicular m2=-1/m1=-1/(3/2)=-2/3 y-yo=m(x-x0) y-3=-2/3(x+6) y=-2/3x-1